OpenStudy (anonymous):

Find an equation of the line containing the given point an parallel to the given line. (-7,-2)and (-3,-5) The equation of the line is Y=

OpenStudy (anonymous):

y = (-3/4)x - (29/4)

OpenStudy (anonymous):

Well you can find your slope by doing (y2-y1)/(x2-x1)... y2 = -5 y1 = -2 x2 = -3 x1 = -7 (-5 - -2)/(-3 - -7) (-3)/(4) Your slope = -3/4

OpenStudy (anonymous):

So since we want our answer in the form of y=mx + b...and since we already have m...we can take one of the two points we were given (it doesn't matter which) and plug it in and solve for b. I'll use (-7, -2)... y=mx+b -2 = (-3/4) * (-7) + b -2 = (21/4) + b Convert -2 into a fraction for the upcoming fractional subtraction... -2 = -8/4 (-8/4) = (21/4) + b Subtract (21/4) from both sides... (-8/4) - (21/4) = (21/4) + b - (21/4) (-29/4) = b

OpenStudy (anonymous):

Excellent example..

OpenStudy (anonymous):

can u help me with one more problem please